to get a B in math, Alex must average 80 on five test. Scores on the first four test were 83, 77, 81 and 78. What is the lowest score that she can get on the last text and still get a B?
Answer by ramkikk66(644)

Average is the total of the scores in all the tests / number of tests.
Therefore, total of scores in all tests = Average * number of tests

To get an average of 80 in 5 tests, she must score a total of 5 * 80 = 400
In the first 4 tests, she has scored a total of (83+77+81+78) = 319
So, to get a total of 400 she must score at least (400 - 319) = 81 in the final 
test. 
Lowest score she should get in the 5th test = 81.
Hope you got it :)